OpenStudy (anonymous):

Use the Addition Formula for Sine to prove the Double-Angle Formula for Sine. sin 2x = 2 sin(x) cos(x) Rewrite 2x as x + x, and use the Addition Formula for Sine to simplify. sin 2x= sin(x + x)

OpenStudy (anonymous):

oh, right. do exactly what it says. replace x + y in the formula by x + x and write what you have, it will give you exactly what you want.

OpenStudy (anonymous):

first write \[\sin(x+y)=\sin(x)\cos(y)+\cos(x)\sin(y)\] then put and x everywhere you see a y and you will get the answer.

OpenStudy (anonymous):

I still don't understand it.

OpenStudy (anonymous):

try what i wrote above and see what happens. it is clear that \[x+x=2x\] right? so if you have a formula for \[\sin(x+y)\] you can make one for \[\sin(x+x)\]

OpenStudy (anonymous):

Ok!

OpenStudy (anonymous):

if you are still confused after you do it write again, but i bet it will work out on the first try

OpenStudy (anonymous):

sin (x+y) = cos x sin x + cos x sin x = cos^2 x + sin^2 x